Volume of a gas at NTP is 1.12 × 10–7 cc. The number of molecules is thus equal to:


        (a)    3.01 × 1012


        (b)    3.01 × 1018


        (c)    3.01 × 1024


        (d)    3.01 × 1030


Grade:11

2 Answers

Kevin Nash
10 years ago

Ans - (a)

Abhishek Sharma
10 years ago

1 mole of gas occupies 22.4 litres at STP
1 mole of gas contains Avogadros number of molecules = 6.022*10^23.
You have not given any units for the volume.
The answer would be (volume/22.4)*Avogadros number = (1.12*10^-7/22.4)*6.022\*10^34
= 0.3011*10^27 if volume is in litres
=0.3011*10^24 = 3.011*10^23 if volume is in m^3.
Answer is c).